OpenStudy (anonymous):

The roof of a tower in a castle is shaped like a cone. The height of the roof is 30 ft, and the radius of the base is 15 ft. What is the area of the roof?

OpenStudy (anonymous):

is this the surface area of the roof or the volume?

OpenStudy (anonymous):

It should be for the surface area I think, which is given by πr(r+(h^2+r^2)^1/2) Which is quite easy really: 30^2 + 15^2 = 1125 1125^1/2 = 33.541 giving π*15(15+33.541) = 15π(48.541) =2287.44